OpenStudy (thisisthepartofever):

I'M SO CONFUSED I WILL MEDAL FAN! does 9 + 10 equal 21 or 19? HELP

OpenStudy (thisisthepartofever):

LIKE EVERYBODY SAYS ITS 21 HELP

OpenStudy (ovidia):

19

OpenStudy (thisisthepartofever):

how do you know @ovidia

OpenStudy (ovidia):

they say 21 off a vine but really 10+9 is 19 dude jajaja

OpenStudy (thehelpneeded):

21 is off of a troll

OpenStudy (solomonzelman):

9 + 10 = \(\rm \small How~many~times~can~people~say~the~same~trivial~joke~over~and~over~again?!\)

OpenStudy (thisisthepartofever):

OHH i think i get it

OpenStudy (mathmale):

"LIKE EVERYBODY SAYS ITS 21 HELP" This is a great reason for NOT listening to "EVERYBODY." I'd suggest you start doing your own thinking and experimentation.

OpenStudy (mathmale):

What does the symbol " + " mean to you?

OpenStudy (solomonzelman):

- | = +

OpenStudy (thisisthepartofever):

well you're adding when you have the + symbol. i see, basically, you have the number 10, and if you add 9 to it, you get 19, so 19 is correct @SolomonZelman @mathmale

OpenStudy (hanamalik):

Was it really that hard to realize the 9 + 10= 19? Count on your fingers?

OpenStudy (solomonzelman):

On the average basis, anyone who knows how to spell "confused" also knows 10+9=19.

OpenStudy (mathmale):

Let's go easy (very easy) on the insults, please.

OpenStudy (solomonzelman):

Why do people say 21, and not 17?

OpenStudy (mathmale):

Ignorance, perhaps. Such a great reason for NOT listening to random people as a way of arriving at answers. Doing your own work would help most.

OpenStudy (mathmale):

You could also try experimentation: Make up your own table: 10+1=11 10+2=12 . . . . 10+8=18 10+9=??